Find an equation of the line with the slope that passes through the given point. Write the equation in the form Ax+By=C
So I got a question that says: m=-3/8, (4,5) I did 5=-3/8(4)+b and ended up with 5=-12/8+b. Is that correct? I tried to find b with that equation but I think I did it wrong. I got -7/2. Am I right or wrong? Can you explain to me how to find b? thank you.

Answer by richwmiller(17219)

m=-3/8, (4,5)
5=-3/8*(4)+b
5=-3/2+b
10/2+3/2=b
13/2=b
6.5 or 6 1/2 or 13/2 =b
y=-3/8x+13/2
but you need it in Ax+By=C form
multiply by 8
y=-3/8x+13/2
3/8x+y=13/2
3x+8y=52